OpenStudy (anonymous):

I need help on this Find the values of the indicated functions. The values are approximate. Given cos=0.0326, find sin and tan.

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

i know i do pythagreon therom

hero (hero):

Use \(\cos^2\theta + \sin^2\theta = 1\) to figure out \(\sin \theta\)

hero (hero):

\(0.0326^2 + \sin^2\theta = 1\) \(\sin^2\theta = 1 - 0.0326^2\)

hero (hero):

\(\sin \theta = \sqrt{1 - (0.0326)^2}\)

OpenStudy (anonymous):

i get 1.000531239 for r

OpenStudy (anonymous):

0.9674

OpenStudy (anonymous):

for r

hero (hero):

You don't need to find an r for this

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so sin is 0.9995 and tan is 30.66